what is the exact value of Tan(105 degrees)
Answer by Alan3354(69443)

what is the exact value of Tan(105 degrees)
-------------------
tan(A+B) = (tanA + tanB)/(1 - tanA*tanB)
tan(105) = tan(45 + 60)
= (1 + sqrt(3))/(1 - sqrt(3))
= (1 + sqrt(3))^2/(-2)
= (1 + 3 + 2sqrt(3))/-2
= -2 - sqrt(3)